OpenStudy (anonymous):

find the area of the kite ? A.) 253.92 m2 B.) 190.44 m2 C.) 380.88 m2 D.) 126.96 m2

OpenStudy (anonymous):

any pictures?

OpenStudy (anonymous):

I got 126.96

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

hmm im not sure how did you got that answer ?

OpenStudy (shamim):

area of any triangle is= 1/2*base*height

OpenStudy (shamim):

we can devide ur given kite into two triangles. one is left triangle nd another is right triangle

OpenStudy (shamim):

area of left triangle is =(1/2)*(9.2+9.2)*6.9

OpenStudy (shamim):

area of right triangle is=(1/2)*(9.2+9.2)*13.8

OpenStudy (shamim):

now u hv to add those two area nd u will get the area of ur given kite